Halloween Forum banner

American DJ Fog Storm 1700 HD stopped working, repair.

13556 Views 22 Replies 6 Participants Last post by  neils
This is just a short post to pass something along (kind of busy today).

My fog storm 1700 HP worked fine yesterday when I tested it. But last night when I did a run through, it had crapped out.

Was not even warm. It appears the low fluid sensor magnetic float stuck, turning on the little LED in the front and turning off the ability to make fog. Fog tank was more than half full. So it just sat there for the evening.

Today when I had a chance to take it apart I found the overtemp switch on the heater assembly had not reset, pushing it's little plunger fixed that. I'm assuming when it sat there most of the night and did not run once it got hot enough at some point to trip the switch, which might be normal and happen often, but we would never see that, if it resets as it should have.

I found a youtube from American DJ where they say a common problem with their 1200 and 1700 that get sent back is the magnetic float gets stuck. They say if you have fluid and still see the green LED on the front panel, bang the tank up and down a couple times and see if it goes out.

For a long time I couldn't get mine to free up, and was going to just jump the switch in the black plastic housing (under tank). I saw it supplies 12Vdc but did not get the chance to determine if it was a reed or Hall switch. Before I got a small rod and poked in down the floats center hole, finally freeing it.

Hope this helps in case someone else has one of these POS, um.. I mean fine machines...
1 - 1 of 23 Posts
When fog began making my throat itch and I couldn't talk, I decided that it was time to quit using any fog at all. (And I never really did over do it , either in my opinion.)
1 - 1 of 23 Posts
This is an older thread, you may not receive a response, and could be reviving an old thread. Please consider creating a new thread.